OpenStudy (anonymous):

Given the equation x^2-6x+y^2-8y-299=0 complete the square for both x and y. What is the result?

OpenStudy (anonymous):

\[(x-3)^2+(y-4)^2=299+3^2+4^2\]

OpenStudy (anonymous):

i get \[(x-3)^2+(y-4)^2=324

OpenStudy (anonymous):

\[(x-3)^2+(y-4)^2=324\]

OpenStudy (anonymous):

alright thanks mate

OpenStudy (anonymous):

gives you a circle with center at (3,4) and radius of 18 (since 18^2=324)

OpenStudy (anonymous):

yw